write the equation of the circle satisfying the given condition: concentric with the circle 4x^2 + 4y^2 - 8x + 4y + 1=0 and tangent to the line y= -3/2.
Answer by AnlytcPhil(1116)

write the equation of the circle satisfying the given condition:
concentric with the circle 4x^2 + 4y^2 - 8x + 4y + 1=0 and tangent to the line
y= -3/2.

First we write the equation of the circle

4x² + 4y² - 8x + 4y + 1 = 0

In the form 

(x - h)² + (y-k)² = r²

4x² + 4y² - 8x + 4y + 1 = 0

Divide thru by 4, the common coefficient of x² and y²

x² + y² - 2x + y + 1/4 = 0

Rearrange so that the x terms and the y terms are together 
and the constant term on the right. 

x² - 2x + y² + y = -1/4 

Place the x terms and the y terms in parentheses

(x² - 2x) + (y² + y) = -1/4

Complete the square in both parentheses by the rule
1. Multiply the coefficient of the second term in each parentheses by 1/2
2. Square the result
3. Add at the end of each parentheses and to the opposite side.

(x² - 2x + 1) + (y² + y + 1/4) = -1/4 + 1 + 1/4

Factor each parenthetical expression as the square of a binomial on the left
and combine the terms on the right:

(x - 1)² + (y + 1/2)² = 1

Compare this to 

(x - h)² + (y - k)² = r²

h=1, k = -1/2, r² = 1, so r = 1

The center = (h,k) = (1, -1/2) abd the radius is 1

The desired circle has the same center (1, -1/2) but a different radius.

It is to be tangent to the line y = -3/2 which is a horizontal line.

The point (1, -3/2) is the point on the line y = -3/2 which is
directly below the center (1, -1/2), thus this point, (1, -3/2) must
be the point of tangency. 

The radius of the desired circle is the distance between the center (1, -1/2)
and the point of tangency (1, -3/2) directly below it.  This distance is
(-1/2) - (-3/2) = -1/2 + 3/2 = 2/2 = 1.

But 1 is the radius of the given circle!!!!!!  So the given circle
4x² + 4y² - 8x + 4y + 1 = 0 is ITSELF tangent to the line y = -3/2. 

Thus there is something wrong with your problem unless your teacher allows
you to say that a circle is concentric with ITSELF!!!

Edwin
AnlytcPhil@aol.com